OpenStudy (anonymous):

A right cylinder has a diameter of 50 mm and a height of 40 mm as shown. What is the volume of the cylinder?

OpenStudy (jdoe0001):

\(\bf \textit{volume of a cylinder}=\pi r^2 h\qquad h=height\qquad r=radius\to \frac{diameter}{2}\)